Graph the following system of equations. y=x+2 y=-x+2
Answer by jim_thompson5910(35256)

Use a graphing calculator to graph y=x+2 (shown in green) y=-x+2 (shown in blue)



The two lines intersect at (0,2)

So the solution to the system of equations is (0,2), basically, x = 0 and y = 2.